Please help. I am so confused with all the equations.


Consider an open economy with S = 0.2Y, M = 0.15Y, T = 0.15Y, where S is the savings function, M is the induced imports and T is the tax function.

Further assume the following autonomous components: C0 = 450, G = 950, I = 500; M = 0; and X = 1 000.


3.1.1 Calculate the open economy multiplier. (3)

3.1.2 Calculate the equilibrium income. (2)

3.1.3 Calculate the value of savings. (3)

3.1.4 Calculate the value of net exports, and comment on the external position of the economy based on your answer.(2)


These are my answers, please let me know if I am on the right track:


3.1 Open economy multiplier = 1 / MPS + MPM + MPT

           = 1 / 0.2 + 0.15 + 0.15 = 1 / 0.5

= 2

 

3.2      Y = C + I + G + (X – M)

           Y = 450 + 500 + 950 + 1000

           Y = 2900

 

3.3     S = 0.2Y

           = 0.2 (2900)

           = 580

 

3.4      Net exports = (X-M) = 1000

 

           The country exports more goods than it imports and therefore has a trade surplus.

 


1
Expert's answer
2021-11-25T10:23:57-0500

3.1) Open economy multiplier "=\frac {1}{MPS + MPM + MPT}"

Where;

MPS is marginal propensity to save

MPM is marginal propensity to import

MPT is marginal propensity to tax

"=\frac {1}{0.2 + 0.15 + 0.15}=0.2"


3.2) Y = C + I + G + (X – M)


Y = 450 + 500 + 950 + 1000


Y = 2900


3.3) S = 0.2Y


"= 0.2 \times2900"


= 580

d)Net exports = (X-M)

Where X represents exports and M represents imports.


"=1000-0=1000"


The country has a trade surplus since it exports more goods than it imports